# On the Low-Temperature MCMC threshold:

### v1

@@ 12.2.1. The Pure Noise Model
$\frac{1}{n}\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\geq\frac{1}{n}\log\binom{p}{k}+\log\tau-\tau^{2}+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
Then by definition of $\tau$, we have $\tau\geq R^{1/2}\exp\left(-\frac{k\log(p/k^{2})}{2n}\right)$, and so
$\frac{1}{n}\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\geq\frac{1}{n}\log\binom{p}{k}+\frac{1}{2}\log R-\frac{k\log\frac{p}{k^{2}}}{2n}+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
We can apply Lemma 12.8 (since $k\leq\sqrt{p})$ to see that $\frac{1}{n}\log\binom{p}{k}-\frac{k}{2n}\log\frac{p}{k^{2}}\geq 0$, which tells us that
$\frac{1}{n}\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\geq\frac{1}{2}\log R+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
@@
$\frac{1}{2}\log\frac{2}{\pi}+\frac{2k\log\frac{p}{k}}{12n}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
@@
$Z_{\tau\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]$
@@
$\mathbb{P}\left(Z_{\tau\sqrt{k},\infty}<\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\right)\leq\mathbb{P}\left(|Z_{\tau\sqrt{k},\infty}-\mathbb{E}\left[Z_{\tau\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]\right)$
@@
$\mathbb{P}\left(|Z_{\tau\sqrt{k},\infty}-\mathbb{E}\left[Z_{\tau\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]|Y\right)$ $\leq 4\min\left[\frac{\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}^{2}|Y\right]}{\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]^{2}}-1,1\right]$
@@
$\mathbb{P}\left(|Z_{\tau\sqrt{k},\infty}-\mathbb{E}\left[Z_{\tau\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]|Y\right)$ $\leq 4\min\left[\frac{\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}^{2}|Y\right]}{\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]^{2}}-1,1\right]$
@@
as $k\rightarrow\infty$ by Proposition 12.11. This proves the claim that $Z_{\tau\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]$ with high probability. We combine this result with Lemma 12.14 to conclude that w.h.p.,
$n^{-1}\log Z_{\tau\sqrt{k},\infty}$ $\geq n^{-1}\log\mathbb{E}\left[Z_{\tau\sqrt{k},\infty}|Y\right]-\frac{\log 2}{n}$
@@
and therefore w.h.p., $Z_{\tau\sqrt{k},\infty}\geq\frac{1}{2}R^{\frac{n}{4}}\geq R^{\frac{n}{5}}$ for sufficiently large $D_{0}$, which concludes the proof of Theorem 12.7. ∎
@@ 12.2.2. The Linear model
By assumption of Theorem 10.1, we have $n\leq\alpha k\log(p/k)/(\log D_{0})$, which therefore gives us
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)+O\left(\frac{k^{2}}{p}\right)}{\log D_{0}}$
Since $k^{2}/p\rightarrow 0$, this means that for large enough $D_{0}$, it holds that
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)}{\log 2D_{0}}$
Using the extra factor of 2 in front of $D_{0}$ lets us replace $\exp(-k\log\left(\frac{p-k}{k}\right)/n)$ by $\exp(-k\log\left(\frac{p}{k}\right)/n)$ in the upper bound on $n^{-\frac{1}{2}}\lVert Y^{\prime}-X^{\prime}v\rVert_{2}$ in Corollary 12.15, which completes the proof of Theorem 10.1. ∎
13. Acknowledgment

### v2

@@ 12.2.1. The Pure Noise Model
$\frac{1}{n}\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\geq\frac{1}{n}\log\binom{p}{k}+\log t_{0}-t_{0}^{2}+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
Then by definition of $t_{0}$, we have $t_{0}\geq R\exp\left(-\frac{k\log(p/k)}{n}\right)$, and so
$\frac{1}{n}\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\geq\frac{1}{n}\log\binom{p}{k}+\log R-\frac{k\log\frac{p}{k}}{n}+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
Since $\binom{p}{k}\geq(p/k)^{k}$, we see that $\frac{1}{n}\log\binom{p}{k}\geq\frac{1}{n}k\log\frac{p}{k}.$ Apply this to see that
$\frac{1}{n}\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\geq\frac{1}{2}\log R+\frac{1}{2}\log\frac{2}{\pi}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
@@
$\frac{1}{2}\log\frac{2}{\pi}+\frac{\alpha k\log\frac{p}{k}}{4n}-\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}^{2}}{k}$
@@
$Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]$
@@
$\mathbb{P}\left(Z_{t_{0}\sqrt{k},\infty}<\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\right)\leq\mathbb{P}\left(|Z_{t_{0}\sqrt{k},\infty}-\mathbb{E}\left[Z_{t_{0}\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]\right)$
@@
$\mathbb{P}\left(|Z_{t_{0}\sqrt{k},\infty}-\mathbb{E}\left[Z_{t_{0}\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]|Y\right)$ $\leq 4\min\left[\frac{\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}^{2}|Y\right]}{\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]^{2}}-1,1\right]$
@@
$\mathbb{P}\left(|Z_{t_{0}\sqrt{k},\infty}-\mathbb{E}\left[Z_{t_{0}\sqrt{k}\infty}|Y\right]|\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]|Y\right)$ $\leq 4\min\left[\frac{\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}^{2}|Y\right]}{\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]^{2}}-1,1\right]$
@@
as $k\rightarrow\infty$ by Proposition 12.11. This proves the claim that $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]$ with high probability. We combine this result with Lemma 12.14 to conclude that w.h.p.,
$n^{-1}\log Z_{t_{0}\sqrt{k},\infty}$ $\geq n^{-1}\log\mathbb{E}\left[Z_{t_{0}\sqrt{k},\infty}|Y\right]-\frac{\log 2}{n}$
@@
and therefore w.h.p., $Z_{t_{0}\sqrt{k},\infty}\geq\frac{1}{2}R^{\frac{n}{4}}\geq R^{\frac{n}{5}}$ for sufficiently large $D_{0}$, which concludes the proof of Theorem 12.7. ∎
@@ 12.2.2. The Linear model
By assumption of Theorem 10.1, we have $n\leq\alpha k\log(p/k)/(3\log D_{0})$, which therefore gives us
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)+O\left(\frac{k^{2}}{p}\right)}{3\log D_{0}}$
Since $k^{2}/p\rightarrow 0$, this means that for large enough $D_{0}$ and large enough values of $k$, it holds that
$n\leq\frac{\alpha k\log\left(\frac{p-k}{k}\right)}{\log 2D_{0}}$ (12.16)
because $3\log D_{0}>\log(2D_{0})$ when $D_{0}>\sqrt{2}$.
We can then apply Corollary 12.15 with $2D_{0}$ in place of $D_{0}$ to get that there are at least $R^{n/5}$ zero-overlap $v$’s that satisfy
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (12.17)
Finally, note that we can write
$\frac{k\log\frac{p-k}{k}}{n}=\frac{k\log\frac{p}{k}}{n}\cdot\frac{\log\frac{p-k}{k}}{\log\frac{p}{k}}$
Note also that $\log(\frac{p-k}{k})/(\log\frac{p}{k})\geq 1-o(1)$ as $k\rightarrow\infty$. It therefore holds that as $k$ grows,
$-\frac{k\log\frac{p-k}{k}}{n}\leq-\frac{1-o(1)}{n}k\log\frac{p}{k}.$
which means that there are at least $R^{n/5}$ zero-overlap $v$’s that satisfy
$n^{-\frac{1}{2}}\lVert Y-Xv\rVert_{2}\leq R\sqrt{2k+\sigma^{2}}\exp\left(-\frac{(1-o(1))k\log\left(\frac{p-k}{k}\right)}{n}\right)$ (12.18)
which completes the proof of Theorem 10.1. ∎
13. Acknowledgments